discuss the function f(x) = x^4/4 - 3x^2/2 for where mimimum and maximum


1
Expert's answer
2021-02-24T06:00:20-0500

Solution.

"f(x)=\frac{x^4}{4}-\frac{3x^2}{2}"

Find "f'(x)=4\frac{x^3}{4}-3\frac{2x}{2}=x^3-3x."

Make and solve equation "f'(x)=0."


"x^3-3x=0,\newline
x(x^2-3)=0,\newline
x_1=0, \text{ or }x_2=\sqrt{3},\text{ or }x_3=-\sqrt{3}."


Investigate the sign of the derivative at each interval:



From here "\sqrt{3}" and "-\sqrt{3}" are minimum points, is maximum point.

"f_{min}(\sqrt{3})=\frac{\sqrt{3}^4}{4}-\frac{3\sqrt{3}^2}{2}=\frac{9}{4}-\frac{9}{2}=
-\frac{9}{4}=-2\frac{1}{4}."

"f_{min}(-\sqrt{3})=\frac{(-\sqrt{3})^4}{4}-\frac{3(-\sqrt{3})^2}{2}=\frac{9}{4}-\frac{9}{2}=
-\frac{9}{4}=-2\frac{1}{4}."

"f_{max}(0)=0."

Answer. "f_{min}=-2\frac{1}{4}, f_{max}=0."